OpenStudy (anonymous):

Create three fractions by replacing the word ‘blank’ in each fraction with a variable or number. You can use the same value or three different replacement values. Explain all the work in simplifying your new addition problem. 3 blank blank -----+ ---- - ----- blank a^2 a^3

OpenStudy (anonymous):

\[\frac{3}{a^{3}} + \frac{2}{a^{2}} - \frac{4}{a^{3}} = \frac{3}{a^{3}} + \frac{2a}{a^{3}} - \frac{4}{a^{3}} = \frac{3 + 2a - 4}{a^{3}} = \frac{2a - 1}{a^{3}}\] I used the values a^3, 2, and 4.

OpenStudy (anonymous):

can you explain what you did please ? i don't understand this at all

OpenStudy (anonymous):

The look of any fraction can be changed by multiplying it by a value of 1, e.g: \[1 = \frac{2}{2} = \frac{2500}{2500}\] Multiply each fraction by 1 in one form or another so that they have the same denominator. Then you add them together and simplify as needed.